# Vulnerability-Aware Alignment: Mitigating Uneven Forgetting in Harmful Fine-Tuning

### v1

@@ Abstract
Figure 1: Overview of the Vulnerability-Aware Alignment. In the first stage, the vulnerability of the alignment dataset is analyzed, and the data is partitioned into two groups using the strategy described in Section 2.2. In the second stage, the adversarial learning framework consists of a LLM and an adversarial sampler, parameterized by a probability simplex over the partitioned groups. At each alignment step, the sampler samples a hard group and applies a group-specific perturbation to challenge the LLM. After training, only the LLM is retained, and the sampler is discarded.
@@ 2.2 Characterizing Data Vulnerability in HFT
We classify alignment examples as vulnerable if ForgotNum $>$ 0, and as invulnerable otherwise. As shown in Figure 1 (left), this grouping is used as prior knowledge in our method. Notably, this procedure is fully data-driven and does not rely on access to the actual downstream fine-tuning distribution, making it broadly applicable across real-world alignment settings.
@@ 3.2 Alignment Training via GDRO
where $q$ is a vector belonging to the $(m-1)$-dimensional probability simplex $\Delta_{m-1}$.
@@ 5 Related Work
Distributionally Robust Optimization. Distributionally robust optimization (DRO) optimizes the objective function over ambiguity sets, often defined as balls centered on the empirical distribution (Ben-Tal et al., 2013; Lam & Zhou, 2015; Duchi et al., 2016; Miyato et al., 2018). Prior applications of DRO have addressed distributional shifts, including covariate shift (Oren et al., 2019; Chen et al., 2025), label shift (Hu et al., 2018), and group shift (Sagawa et al., 2020). To the best of our knowledge, we are the first to apply DRO to defend against the harmful fine-tuning problem by encouraging LLMs to learn equally well and robustly on both vulnerable and non-vulnerable examples.
@@ 6 Conclusion
Our work highlights the potential of mitigating harmful fine-tuning from a data perspective. By focusing on the inherent vulnerability of data, we have shown that addressing data-dependent vulnerabilities can significantly improve the safety and reliability of large language models. Future research could explore more nuanced ways to categorize data, such as utilizing a continuous spectrum of vulnerability rather than a binary classification of vulnerable versus invulnerable. Additionally, our method is orthogonal to other alignment-stage methods, and combining VAA with these existing methods could offer even more robust protections against harmful fine-tuning. Ultimately, our work paves the way for more comprehensive and robust solutions to alignment breakdowns in the future.

### v2

@@ Abstract
Figure 1: Overview of the Vulnerability-Aware Alignment. In the first stage, the vulnerability of the alignment dataset is analyzed, and the data is partitioned into two groups using the strategy described in Section $2.$ In the second stage, the adversarial learning framework consists of a LLM and an adversarial sampler, parameterized by a probability simplex over the partitioned groups. At each alignment step, the sampler samples a hard group and applies a group-specific perturbation to challenge the LLM. After training, only the LLM is retained, and the sampler is discarded.
@@ 2.2 Characterizing Data Vulnerability in HFT
We classify alignment examples as vulnerable if ForgotNum >> 0, and as invulnerable otherwise. As shown in Figure 1 (left), this grouping is used as prior knowledge in our method. Notably, this procedure is fully data-driven and does not rely on access to the actual downstream fine-tuning distribution, making it broadly applicable across real-world alignment settings.
@@ 3.2 Alignment Training via GDRO
where $q$ is a vector belonging to the (m1)(m-1)-dimensional probability simplex $\Delta_{m-1}$.
@@ 5 Related Work
Distributionally Robust Optimization. Distributionally robust optimization (DRO) optimizes the objective function over ambiguity sets, often defined as balls centered on the empirical distribution (Ben-Tal et al., 2013; Lam & Zhou, 2015; Duchi et al., 2016; Miyato et al., 2018). Prior applications of DRO have addressed distributional shifts, including covariate shift (Oren et al., 2019; Chen et al., 2024b, 2025), label shift (Hu et al., 2018), and group shift (Sagawa et al., 2020). To the best of our knowledge, we are the first to apply DRO to defend against the harmful fine-tuning problem by encouraging LLMs to learn equally well on both vulnerable and non-vulnerable examples.
@@ 6 Conclusion
Our work highlights the potential of mitigating harmful fine-tuning from a data-centric perspective. By focusing on the inherent vulnerability of data, we have shown that addressing data-dependent vulnerabilities can significantly improve the safety and reliability of LLMs. Notably, VAA is orthogonal to existing alignment-stage methods, and combining it with such techniques may yield even more robust defenses against harmful fine-tuning. This work lays the groundwork for more comprehensive and resilient solutions to alignment breakdowns during fine-tuning.
Limitations The current data partitioning strategy is relatively simple and relies on a pseudo fine-tuning process. Future research could explore more nuanced ways of categorizing data, such as employing a continuous spectrum of vulnerability—e.g., based on uncertainty estimation (Gawlikowski et al., 2022) or factuality metrics (Lin et al., 2022; Chen et al., 2023)—without the need for additional pseudo fine-tuning stages. Moreover, while VAA is effective in reducing alignment breakdowns during customized fine-tuning, it does not prevent them entirely. Therefore, integration with complementary techniques such as AI watermarking (Kirchenbauer et al., 2023; Chen et al., 2024a) may be necessary to achieve more reliable protection and traceability for open-source LLMs.